# Tokenizing Single-Channel EEG with Time-Frequency Motif Learning

### v3

@@ Abstract
Foundation models are reshaping EEG analysis, yet an important problem of EEG tokenization remains a challenge. This paper presents TFM-Tokenizer, a novel tokenization framework that learns a vocabulary of time-frequency motifs from single-channel EEG signals and encodes them into discrete tokens. We propose a dual-path architecture with time–frequency masking to capture robust motif representations, and it is model-agnostic, supporting both lightweight transformers and existing foundation models for downstream tasks. Our study demonstrates three key benefits: Accuracy: Experiments on four diverse EEG benchmarks demonstrate consistent performance gains across both single- and multi-dataset pretraining settings, achieving up to $17\%$ improvement in Cohen’s Kappa over strong baselines. Generalization: Moreover, as a plug-and-play component, it consistently boosts the performance of diverse foundation models, including BIOT and LaBraM. Scalability: By operating at the single-channel level rather than relying on the strict 10–20 EEG system, our method has the potential to be device-agnostic. Experiments on ear-EEG sleep staging, which differs from the pretraining data in signal format, channel configuration, recording device, and task, show that our tokenizer outperforms baselines by $14\%$. A comprehensive token analysis reveals strong class-discriminative, frequency-aware, and consistent structure, enabling improved representation quality and interpretability. Code is available at https://github.com/Jathurshan0330/TFM-Tokenizer.
@@ 1 Introduction
Various studies have shown that developing an effective tokenization is a non-trivial task in general, as it is influenced by multiple factors (Schmidt et al., 2024). In this paper, we recognize and focus on three key challenges of EEG tokenization. 1) Tokenization target: real-world EEG recordings exhibit diverse formats due to varying devices, channel configurations, and recording lengths (Yang et al., 2024). We argue that tokenizers should be trained and operated at the single-channel level to learn channel-agnostic discrete tokens. This design enables flexible adaptation to multi-channel tasks and can generalize to non-standard EEG devices. In Section 4.4, we provide scalability experiments on ear-EEG settings. 2) Token resolution: in NLP, tokenization can be defined at different resolutions (characters, subwords, words), each reflecting different assumptions about semantic granularity. However, EEG signals are characterized by diverse oscillatory (e.g., alpha, beta) (Pradeepkumar et al., 2024) and transient patterns (e.g., spikes) (Chen et al., 2022). Thus, effective tokens must represent such underlying motifs (Xu et al., 2023) that reflect distinct neural or physiological events. However, these motifs are often distorted by noise, amplitude scaling, and temporal warping, making it challenging to design robust EEG tokenization methods. 3) Tokenization learning objective: EEGs exhibit various temporal variations, manifested as a mixture of low- and high-frequency components that co-occur and are intermixed in complex ways. Relying solely on capturing time‑based motifs into discrete tokens risks losing important spectral structure. We therefore argue that the tokenization learning objective should incorporate time–frequency representations, enabling tokens to encode more meaningful EEG motifs.
@@
  • Broad Evaluation across Foundation Models and Devices. Extensive experiments across four datasets show that our method outperforms strong baselines, achieving up to a $17\%$ gain over the baseline model on TUEV dataset. We also evaluate cross-device scalability on an ear-EEG sleep staging task, using electrodes outside the standard 10–20 EEG system, where our tokenizer outperforms baselines by $14\%$. Beyond performance, we comprehensively analyze token quality, including token consistency, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
@@ 4.2 How Does TFM-Tokenizer Compare to Existing Baselines?
Table 1 reports results on TUEV (event classification) and TUAB (abnormal detection), while Table 2 summarizes performance on IIIC-Seizure (seizure type classification) and CHB-MIT (seizure detection). Our TFM-Tokenizer paired with a downstream transformer consistently outperforms all baselines in both experiment settings. On the challenging six-class event-type classification task in TUEV, it achieves a $5\%$ gain in Cohen’s Kappa in the single-dataset setting and a notable $17\%$ improvement ($0.5273\rightarrow 0.6189$) in the multi-dataset setting over the next best baseline. On IIIC-Seizure, which is another six-class classification task, TFM-Tokenizer improves Cohen’s Kappa by $36\%$ over the next best baseline LaBraM ($0.3658\rightarrow 0.4979$, p = 1.5e-4) in multiple dataset settings, demonstrating the strong capability of our tokenizer in modeling class-discriminative features for complex clinical EEG tasks. Additionally, it is worth noting that TFM-Tokenizer achieves better performance with fewer parameters, being 3 times smaller than LaBraM and 1.5 times smaller than BIOT. The ability to achieve best performance with low model size can be attributed to our tokenization approach, which compresses the EEG into a token sequence, thereby reducing data complexity. Notably, the TFM-Tokenizer is paired with a lightweight transformer comprising only $\sim$0.7M parameters.
@@
[…]
@@
[…]
@@ 4.4 Does TFM-Tokenizer Scale to Other Brain-signal Types / Devices?
[…]
TFM-Tokenizer $\dagger$ 0.4148 $\pm$ 0.0209 0.3883 $\pm$ 0.0233 0.5174 $\pm$ 0.0141
@@ 6 Reproducibility statement
To support the reproducibility of our work, we provide our complete source code and pretrained model weights at https://github.com/Jathurshan0330/TFM-Tokenizer. The repository includes scripts for data preprocessing, loading, and model training to reproduce our results presented in this paper. In the main text, Section 4.1 outlines our experimental setup, including descriptions of the dataset and baselines. Additional implementation details, such as dataset statistics, preprocessing steps, ear-EEG-specific processing, evaluation metrics, and baseline configurations, are provided in Appendix B.1B.2B.3B.4, and B.5. The Appendix also includes extended experiments across multiple datasets, including frequency learning analysis (Appendix C.1), cross-dataset generalization studies (Appendix C.3), additional results on improving foundation models (Appendix C.4), and further ablation studies. We have made every effort to ensure that our work can be easily reproduced by the community.
@@ B.1 Dataset Statistics and Splits
TUAB $2,339$ $409,455$ $10$ Abnoral EEG Detection
@@ C.3 Token Generalization Assessment through Cross-Dataset Experiments
To evaluate the robustness of our tokenizer, we conducted cross-dataset experiments under two settings: (1) fixing the tokenizer and performing masked token prediction (MTP) & finetuning on a different target dataset and (2) fixing the tokenizer and MTP, followed by finetuning TFM-Encoder only on the target dataset. Results are presented in Table 8, which demonstrates strong generalizability, with our TFM-Tokenizer achieving the best performance on TUEV when pretrained on CHBMIT—outperforming the best-reported result in four dataset settings. These findings highlight the potential of our tokenizer as a foundation for a scalable, universal EEG tokenizer.
@@ C.9 Ablation on Masking
Frequency Bin Masking $0.4673\pm 0.0540$ $0.5193\pm 0.0243$ $0.7536\pm 0.0125$
Frequency Bin $\mathbf{0.4946\pm 0.0392}$ $0.5045\pm 0.0221$ $0.7462\pm 0.0116$
@@
Frequency Bin $0.4943\pm 0.0516$ $\mathbf{0.5337\pm 0.0306}$ $\mathbf{0.7570\pm 0.0163}$
@@
We conducted an ablation study on masking strategies during TFM-Tokenizer pretraining to assess their impact on performance. Results shown in Table 12 indicate that random masking on the spectrogram $S$ performs poorly compared to other strategies, underscoring the need for effective masking to capture frequency and temporal features from EEG. Frequency bin masking significantly improves performance over random masking, with an $8\%$ increase in Cohen’s Kappa ($0.4772\rightarrow 0.5193$) and a $7\%$ increase in balanced accuracy ($0.4351\rightarrow 0.4673$), highlighting the importance of modeling frequency band dynamics. The addition of temporal masking further boosts balanced accuracy by $5\%$ ($0.4673\rightarrow 0.4946$), underscoring the importance of joint temporal-frequency modeling. However, temporal masking results in a decline in Cohen’s Kappa and Weighted F1, which is then resolved by introducing symmetric masking, achieving the overall best performance.

### v4

@@ Abstract
Foundation models are reshaping EEG analysis, yet an important problem of EEG tokenization remains a challenge. This paper presents TFM-Tokenizer, a novel tokenization framework that learns a vocabulary of time-frequency motifs from single-channel EEG signals and encodes them into discrete tokens. We propose a dual-path architecture with time–frequency masking to capture robust motif representations, and it is model-agnostic, supporting both lightweight transformers and existing foundation models for downstream tasks. Our study demonstrates three key benefits: Accuracy: Experiments on four diverse EEG benchmarks demonstrate consistent performance gains across both single- and multi-dataset pretraining settings, achieving up to $11\%$ improvement in Cohen’s Kappa over strong baselines. Generalization: Moreover, as a plug-and-play component, it consistently boosts the performance of diverse foundation models, including BIOT and LaBraM. Scalability: By operating at the single-channel level rather than relying on the strict 10–20 EEG system, our method has the potential to be device-agnostic. Experiments on ear-EEG sleep staging, which differs from the pretraining data in signal format, channel configuration, recording device, and task, show that our tokenizer outperforms baselines by $14\%$. A comprehensive token analysis reveals strong class-discriminative, frequency-aware, and consistent structure, enabling improved representation quality and interpretability. Code is available at https://github.com/Jathurshan0330/TFM-Tokenizer.
@@ 1 Introduction
Various studies have shown that developing an effective tokenization is a non-trivial task in general, as it is influenced by multiple factors (Schmidt et al., 2024). In this paper, we recognize and focus on three key challenges of EEG tokenization. 1) Tokenization target: real-world EEG recordings exhibit diverse formats due to varying devices, channel configurations, and recording lengths (Yang et al., 2024). We argue that tokenizers should be trained and operated at the single-channel level to learn channel-agnostic discrete tokens. This design enables flexible adaptation to multi-channel tasks and can generalize to non-standard EEG devices. In Section 4.4, we provide scalability experiments on ear-EEG settings. 2) Token resolution: in NLP, tokenization can be defined at different resolutions (characters, subwords, words), each reflecting different assumptions about semantic granularity. However, EEG signals are characterized by diverse oscillatory (e.g., alpha, beta) (Pradeepkumar et al., 2024) and transient patterns (e.g., spikes) (Chen et al., 2022). Thus, effective tokens must represent such underlying motifs (Xu et al., 2023) that reflect distinct neural or physiological events. Motifs can be understood as short, recurring patterns in a time series that exhibit limited variability and often carry discriminative significance (Xu et al., 2023). However, these motifs are often distorted by noise, amplitude scaling, and temporal warping, making it challenging to design robust EEG tokenization methods. 3) Tokenization learning objective: EEGs exhibit various temporal variations, manifested as a mixture of low- and high-frequency components that co-occur and are intermixed in complex ways. Relying solely on capturing time‑based motifs into discrete tokens and expecting the model to implicitly infer spectral structure from raw signals risks overlooking important frequency information. We therefore argue that the tokenization learning objective should explicitly incorporate time–frequency representations, enabling the tokenizer to capture band-specific and cross-frequency patterns and to encode more meaningful neural motifs
@@
  • Broad Evaluation across Foundation Models and Devices. Extensive experiments across four datasets show that our method outperforms strong baselines, achieving up to a $11\%$ gain over the baseline model on TUEV dataset. We also evaluate cross-device scalability on an ear-EEG sleep staging task, using electrodes outside the standard 10–20 EEG system, where our tokenizer outperforms baselines by $14\%$. Beyond performance, we comprehensively analyze token quality, including token consistency, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
@@ 4.2 How Does TFM-Tokenizer Compare to Existing Baselines?
Table 1 reports results on TUEV (event classification) and TUAB (abnormal detection), while Table 2 summarizes performance on IIIC-Seizure (seizure type classification) and CHB-MIT (seizure detection). Our TFM-Tokenizer paired with a downstream transformer outperforms the baselines in both experiment settings. On the challenging six-class event-type classification task in TUEV, it achieves a $5\%$ gain in Cohen’s Kappa in the single-dataset setting and a notable $\sim 11\%$ improvement ($0.5588\rightarrow 0.6189$) in the multi-dataset setting over the next best baseline. On IIIC-Seizure, which is another six-class classification task, TFM-Tokenizer improves Cohen’s Kappa by $36\%$ over the LaBraM ($0.3658\rightarrow 0.4979$) and $3\%$ improvement over CBraMod ($0.4792\rightarrow 0.4979$) in multiple dataset settings, demonstrating the strong capability of our tokenizer in modeling class-discriminative features for complex clinical EEG tasks. Additionally, it is worth noting that TFM-Tokenizer achieves better performance with fewer parameters, being 3 times smaller than LaBraM and 1.5 times smaller than BIOT. The ability to achieve best performance with low model size can be attributed to our tokenization approach, which compresses the EEG into a token sequence, thereby reducing data complexity. Notably, the TFM-Tokenizer is paired with a lightweight transformer comprising only $\sim$0.7M parameters.
@@
Models Model TUEV (event type classification) TUAB (abnormal detection) Size Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC Single Dataset Setting SPaRCNet (Jing et al., 2023) 0.79M 0.4161 $\pm$ 0.0262 0.4233 $\pm$ 0.0181 0.7024 $\pm$ 0.0104 0.7896 $\pm$ 0.0018 0.8414 $\pm$ 0.0018 0.8676 $\pm$ 0.0012 ContraWR (Yang et al., 2023) 1.6M 0.4384 $\pm$ 0.0349 0.3912 $\pm$ 0.0237 0.6893 $\pm$ 0.0136 0.7746 $\pm$ 0.0041 0.8421 $\pm$ 0.0104 0.8456 $\pm$ 0.0074 CNN-Transformer (Peh et al., 2022) 3.2M 0.4087 $\pm$ 0.0161 0.3815 $\pm$ 0.0134 0.6854 $\pm$ 0.0293 0.7777 $\pm$ 0.0022 0.8433 $\pm$ 0.0039 0.8461 $\pm$ 0.0013 FFCL (Li et al., 2022) 2.4M 0.3979 $\pm$ 0.0104 0.3732 $\pm$ 0.0188 0.6783 $\pm$ 0.0120 0.7848 $\pm$ 0.0038 0.8448 $\pm$ 0.0065 0.8569 $\pm$ 0.0051 ST-Transformer (Song et al., 2021) 3.5M 0.3984 $\pm$ 0.0228 0.3765 $\pm$ 0.0306 0.6823 $\pm$ 0.0190 0.7966 $\pm$ 0.0023 0.8521 $\pm$ 0.0026 0.8707 $\pm$ 0.0019 Vanilla BIOT (Yang et al., 2024) 3.2M 0.4682 $\pm$ 0.0125 0.4482 $\pm$ 0.0285 0.7085 $\pm$ 0.0184 0.7925 $\pm$ 0.0035 0.8707 $\pm$ 0.0087 0.8691 $\pm$ 0.0033 BIOT (Yang et al., 2024) 3.2M 0.4679 $\pm$ 0.0354 0.4890 $\pm$ 0.0407 0.7352 $\pm$ 0.0236 0.7955 $\pm$ 0.0047 0.8819 $\pm$ 0.0046 0.8834 $\pm$ 0.0041 LaBraM-Base (Jiang et al., 2024b) 5.8M 0.4682 $\pm$ 0.0856 0.5067 $\pm$ 0.0413 0.7466 $\pm$ 0.0202 0.7720 $\pm$ 0.0046 0.8498 $\pm$ 0.0036 0.8534 $\pm$ 0.0027 TFM-Tokenizer (Ours) 1.9M 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163 0.8152 $\pm$ 0.0014 0.8946 $\pm$ 0.0008 0.8897 $\pm$ 0.0008 With Multiple Dataset Pretraining BIOT (Yang et al., 2024) 3.2M 0.5281 $\pm$ 0.0225 0.5273 $\pm$ 0.0249 0.7492 $\pm$ 0.0082 0.7959 $\pm$ 0.0057 0.8792 $\pm$ 0.0023 0.8815 $\pm$ 0.0043 EEGPT (Wang et al., 2024a) 4.7M 0.5670 $\pm$ 0.0066 0.5085 $\pm$ 0.0173 0.7535 $\pm$ 0.0097 0.7959 $\pm$ 0.0021 - 0.8716 $\pm$ 0.0041 NeuroLM-B (Jiang et al., 2024a) 254M 0.4560 $\pm$ 0.0048 0.4285 $\pm$ 0.0048 0.7153 $\pm$ 0.0028 0.7826 $\pm$ 0.0065 0.6975 $\pm$ 0.0081 0.7816 $\pm$ 0.0079 LaBraM-Base (Jiang et al., 2024b) 5.8M 0.5550 $\pm$ 0.0403 0.5175 $\pm$ 0.0339 0.7450 $\pm$ 0.0194 0.7735 $\pm$ 0.0030 0.8531 $\pm$ 0.0028 0.8557 $\pm$ 0.0027 CBraMod (Wang et al., 2024d) 4M 0.5696 $\pm$ 0.0221 0.5588 $\pm$ 0.0273 0.7702 $\pm$ 0.0137 0.5000 $\pm$ 0.0000 0.4938 $\pm$ 0.0443 0.5281 $\pm$ 0.0409 TFM-Tokenizer (Ours) 1.9M 0.5974 $\pm$ 0.0079 0.6189 $\pm$ 0.0302 0.8010 $\pm$ 0.0161 0.8032 $\pm$ 0.0035 0.8886 $\pm$ 0.0032 0.8870 $\pm$ 0.0022
@@
Models Model IIIC Seizure (seizure type classification) CHB-MIT (seizure detection) Size Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC Single Dataset Setting SPaRCNet (Jing et al., 2023) 0.79M 0.5011 $\pm$ 0.0286 0.4115 $\pm$ 0.0297 0.4996 $\pm$ 0.0262 0.5876 $\pm$ 0.0191 0.1247 $\pm$ 0.0119 0.8143 $\pm$ 0.0148 ContraWR (Yang et al., 2023) 1.6M 0.5421 $\pm$ 0.0123 0.4549 $\pm$ 0.0166 0.5387 $\pm$ 0.0138 0.6344 $\pm$ 0.0002 0.2264 $\pm$ 0.0174 0.8097 $\pm$ 0.0114 CNN-Transformer (Peh et al., 2022) 3.2M 0.5395 $\pm$ 0.0144 0.4500 $\pm$ 0.0165 0.5413 $\pm$ 0.0176 0.6389 $\pm$ 0.0067 0.2479 $\pm$ 0.0227 0.8662 $\pm$ 0.0082 FFCL (Li et al., 2022) 2.4M 0.5309 $\pm$ 0.0217 0.4412 $\pm$ 0.0253 0.5315 $\pm$ 0.0277 0.6262 $\pm$ 0.0104 0.2049 $\pm$ 0.0346 0.8271 $\pm$ 0.0051 ST-Transformer (Song et al., 2021) 3.5M 0.5093 $\pm$ 0.0122 0.4217 $\pm$ 0.0151 0.5217 $\pm$ 0.0110 0.5915 $\pm$ 0.0195 0.1422 $\pm$ 0.0094 0.8237 $\pm$ 0.0491 Vanilla BIOT (Yang et al., 2024) 3.2M 0.5762 $\pm$ 0.0034 0.4932 $\pm$ 0.0046 0.5773 $\pm$ 0.0031 0.6640 $\pm$ 0.0037 0.2573 $\pm$ 0.0088 0.8646 $\pm$ 0.0030 BIOT (Yang et al., 2024) 3.2M 0.4458 $\pm$ 0.0183 0.3418 $\pm$ 0.0228 0.4511 $\pm$ 0.0207 0.6582 $\pm$ 0.0896 0.3127 $\pm$ 0.0890 0.8456 $\pm$ 0.0333 LaBraM-Base (Jiang et al., 2024b) 5.8M 0.4736 $\pm$ 0.0101 0.3716 $\pm$ 0.0128 0.4765 $\pm$ 0.0097 0.5035 $\pm$ 0.0078 0.0959 $\pm$ 0.0742 0.6624 $\pm$ 0.1050 TFM-Tokenizer (Ours) 1.9M 0.5775 $\pm$ 0.0042 0.4985 $\pm$ 0.0039 0.5847 $\pm$ 0.0050 0.6750 $\pm$ 0.0392 0.3379 $\pm$ 0.0515 0.8839 $\pm$ 0.0173 With Multiple Dataset Pretraining BIOT (Yang et al., 2024) 3.2M 0.4414 $\pm$ 0.0035 0.3362 $\pm$ 0.0040 0.4483 $\pm$ 0.0033 0.7068 $\pm$ 0.0457 0.3277 $\pm$ 0.0460 0.8761 $\pm$ 0.0284 EEGPT (Wang et al., 2024a) 4.7M 0.4545 $\pm$ 0.0193 0.3502 $\pm$ 0.0255 0.4559 $\pm$ 0.0311 0.6644 $\pm$ 0.0227 0.3373 $\pm$ 0.0264 0.8185 $\pm$ 0.0252 LaBraM-Base (Jiang et al., 2024b) 5.8M 0.4736 $\pm$ 0.0037 0.3658 $\pm$ 0.0033 0.4708 $\pm$ 0.0015 0.5260 $\pm$ 0.0369 0.2138 $\pm$ 0.0523 0.7750 $\pm$ 0.0540 CBraMod (Wang et al., 2024d) 4M 0.5566 $\pm$ 0.0126 0.4792 $\pm$ 0.0167 0.5743 $\pm$ 0.0138 0.6646 $\pm$ 0.0598 0.3469 $\pm$ 0.0281 0.9071 $\pm$ 0.0199 TFM-Tokenizer (Ours) 1.9M 0.5747 $\pm$ 0.0022 0.4979 $\pm$ 0.0038 0.5797 $\pm$ 0.0017 0.6471 $\pm$ 0.0145 0.3554 $\pm$ 0.0264 0.8818 $\pm$ 0.0117
@@ 4.4 Does TFM-Tokenizer Scale to Other Brain-signal Types / Devices?
BIOT-TFM 0.3952 $\pm$ 0.0170 $\uparrow$ 0.3603 $\pm$ 0.0252 $\uparrow$ 0.5033 $\pm$ 0.0165 $\uparrow$
LaBraM-Base 0.3890 $\pm$ 0.0182 0.3322 $\pm$ 0.0232 0.4827 $\pm$ 0.0157
LaBraM-TFM 0.4004 $\pm$ 0.0086 $\uparrow$ 0.3475 $\pm$ 0.0128 $\uparrow$ 0.4864 $\pm$ 0.0118 $\uparrow$
TFM-Tokenizer 0.4148 $\pm$ 0.0209 0.3883 $\pm$ 0.0233 0.5174 $\pm$ 0.0141
@@ 6 Reproducibility statement
To support the reproducibility of our work, we provide our complete source code and pretrained model weights at https://github.com/Jathurshan0330/TFM-Tokenizer The repository includes scripts for data preprocessing, loading, and model training to reproduce our results presented in this paper. In the main text, Section 4.1 outlines our experimental setup, including descriptions of the dataset and baselines. Additional implementation details, such as dataset statistics, preprocessing steps, ear-EEG-specific processing, evaluation metrics, and baseline configurations, are provided in Appendix B.1B.2B.3B.4, and B.5. The Appendix also includes extended experiments across multiple datasets, including frequency learning analysis (Appendix C.1), cross-dataset generalization studies (Appendix C.3), additional results on improving foundation models (Appendix C.4), and further ablation studies. We have made every effort to ensure that our work can be easily reproduced by the community.
@@ B.1 Dataset Statistics and Splits
TUAB $2,339$ $409,455$ $10$ Abnormal EEG Detection
@@ C.3 Token Generalization Assessment through Cross-Dataset Experiments
To evaluate the robustness of our tokenizer, we conducted cross-dataset experiments under two settings: (1) fixing the tokenizer and performing masked token prediction (MTP) & finetuning on a different target dataset and (2) fixing the tokenizer and MTP, followed by finetuning downstream transformer only on the target dataset. Results are presented in Table 8, which demonstrates strong generalizability, with our TFM-Tokenizer achieving the best performance on TUEV when pretrained on CHBMIT—outperforming the best-reported result in four dataset settings. These findings highlight the potential of our tokenizer as a foundation for a scalable, universal EEG tokenizer.
@@ C.9 Ablation on Masking
Frequency Band Masking $0.4673\pm 0.0540$ $0.5193\pm 0.0243$ $0.7536\pm 0.0125$
Frequency Band $\mathbf{0.4946}\pm 0.0392$ $0.5045\pm 0.0221$ $0.7462\pm 0.0116$
@@
Frequency Band $0.4943\pm 0.0516$ $\mathbf{0.5337}\pm 0.0306$ $\mathbf{0.7570}\pm 0.0163$
@@
We conducted an ablation study on masking strategies during TFM-Tokenizer pretraining to assess their impact on performance. Results shown in Table 12 indicate that random masking on the spectrogram $S$ performs poorly compared to other strategies, underscoring the need for effective masking to capture frequency and temporal features from EEG. Frequency band masking significantly improves performance over random masking, with an $8\%$ increase in Cohen’s Kappa ($0.4772\rightarrow 0.5193$) and a $7\%$ increase in balanced accuracy ($0.4351\rightarrow 0.4673$), highlighting the importance of modeling frequency band dynamics. The addition of temporal masking further boosts balanced accuracy by $5\%$ ($0.4673\rightarrow 0.4946$), underscoring the importance of joint temporal-frequency modeling. However, temporal masking results in a decline in Cohen’s Kappa and Weighted F1, which is then resolved by introducing symmetric masking, achieving the overall best performance.
C.10 Masking Ratio Ablation
Table 13: Ablation on frequency band masking ratio used for the pretraining of TFM-Tokenizer on TUEV, IIIC Seizure and CHB-MIT Datasets.
Dataset Frequency Mask Ratio Balanced Acc. Cohen’s Kappa Weighted F1
TUEV 0.5 0.4946 $\pm$ 0.0392 0.5045 $\pm$ 0.0221 0.7462 $\pm$ 0.0116
0.3 0.4306 $\pm$ 0.0187 0.5025 $\pm$ 0.0193 0.7432 $\pm$ 0.0090
0.1 0.3859 $\pm$ 0.0580 0.4308 $\pm$ 0.0755 0.7057 $\pm$ 0.0376
IIIC 0.5 0.5315 $\pm$ 0.0102 0.4427 $\pm$ 0.0143 0.5369 $\pm$ 0.0114
0.3 0.5148 $\pm$ 0.0158 0.4250 $\pm$ 0.0193 0.5222 $\pm$ 0.0167
0.1 0.4381 $\pm$ 0.0032 0.3286 $\pm$ 0.0046 0.4420 $\pm$ 0.0047
Balanced Acc. AUC-PR AUROC
CHB-MIT 0.5 0.6809 $\pm$ 0.0380 0.3335 $\pm$ 0.0182 0.8859 $\pm$ 0.0137
0.3 0.6313 $\pm$ 0.0599 0.3233 $\pm$ 0.0337 0.8708 $\pm$ 0.0187
0.1 0.6530 $\pm$ 0.0486 0.3502 $\pm$ 0.0441 0.8742 $\pm$ 0.0116
We conducted an ablation study to examine how varying the frequency band masking ratio affects model performance and generalization across datasets. All experiments were performed under the single-channel setting, with the temporal masking ratio fixed at 0.5 without symmetric masking, and the results are summarized in Table 13. For the TUEV and IIIC Seizure datasets, a frequency mask ratio of 0.5 yielded the best overall performance. A similar trend was observed in the CHB-MIT dataset, except for Cohen’s Kappa, which showed a slightly higher score at a masking ratio of 0.1. Considering these results along with the added benefit that a 0.5 masking ratio enables more effective use of symmetric masking as a data-augmentation strategy, we selected a frequency mask ratio of 0.5 for our final approach.
C.11 Window Length ($L$) and Hop Size ($H$) Ablation
Table 14: Ablation on window length ($L$) and stride or hop size ($H$) used to segment raw signals and compute STFT for the pretraining of TFM-Tokenizer on TUEV Dataset.
Window Hop Size Balanced Acc. Cohen’s Kappa Weighted F1
Length (s) (s)
0.5 0.25 0.5038 $\pm$ 0.0561 0.6059 $\pm$ 0.0170 0.7935 $\pm$ 0.0112
1.0 0.25 0.4796 $\pm$ 0.0598 0.5761 $\pm$ 0.0171 0.7780 $\pm$ 0.0098
[…]
1.0 0.75 0.4068 $\pm$ 0.0182 0.4868 $\pm$ 0.0210 0.7327 $\pm$ 0.0085
2.0 0.5 0.1726 $\pm$ 0.0093 0.0168 $\pm$ 0.0137 0.5202 $\pm$ 0.0074
2.0 1.0 0.2123 $\pm$ 0.0143 0.1504 $\pm$ 0.0146 0.5748 $\pm$ 0.0087
2.0 1.5 0.3948 $\pm$ 0.0287 0.4042 $\pm$ 0.0282 0.6878 $\pm$ 0.0167
To investigate how window length and stride affect the tokenizer’s ability to capture time–frequency motifs and performance, we conducted an ablation varying both parameters and adjusted the STFT configuration to preserve one-to-one alignment between time and frequency windows. The results, summarized in Table 14, indicate that smaller windows with greater overlap yield the strongest performance. This suggests that shorter segments allow the tokenizer to capture finer-grained motifs that may be lost when using larger windows. For consistency with baselines and prior work, however, we adopt a 1-second window length with a 0.5-second hop size in all reported experiments.
C.12 Token Embedding Size Ablation
Table 15: Token embedding size ablation on TUEV Dataset.
Embedding Balanced Acc. Cohen’s Kappa Weighted F1
Dimension
32 0.4213 $\pm$ 0.0529 0.4974 $\pm$ 0.0165 0.7417 $\pm$ 0.0081
64 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163
128 0.3199 $\pm$ 0.0193 0.1909 $\pm$ 0.0245 0.5700 $\pm$ 0.0276
256 0.3864 $\pm$ 0.0082 0.3575 $\pm$ 0.0157 0.6682 $\pm$ 0.0091
Table 15 summarizes the results of the token embedding size ablation. Performance improves up to an embedding dimension of 64, after which it begins to decline. We do not observe a consistent trend as the embedding size increases, which may be attributed to training instability when using larger embedding dimensions.